The reaction goes to completion, but in the process of washing and drying the precipitate, some was lost. The percent yield for the reaction is 79.9%

79.9%. How many grams of precipitate are recovered?

precipitate recovered:

How many grams of the excess reactant remain? Assume the reaction goes to completion.

excess reactant remaining:


1
Expert's answer
2022-03-20T06:38:15-0400

2KClO3→2KCl+3O2

From the reaction,it can be seen that;


2mol of KClO3= 3mol of O2

1 mol of KClO3=3/2 mol O2

=1.5 mol O2

The yield of the reaction is given in % i.e 79.9%


1 mil KClO3 generates 1.5 mol O2

Since the yield is 79.9%,the number of moles of O2 generated per mole of KClO3 is


1.5mol×79.9/100=1.1985mol


Number of moles ofO2=42.0g/32.0g


=1.31 mol.


Number of moles of KClO3

=1.31mol O2/1.1985 mol O2/KClO3


=1.093 mol KClO3

Weight of KClO3=1.35 mol KClO3×molecular weight of KClO3

=1.093×112.6g


=123.1g